# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ 4. Differentially Private Measurement Design
Given the constraint that each user can contribute at most $r_{i}$ conversions on day $i$ ($1\leq i\leq n$), the user-level neighboring dataset with bounded per-day contributions, denoted as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$, defines a pair of datasets $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})$ as belonging to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ if $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can be obtained from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by substituting all records (with a maximum of $r_{i}$ records on day $i$) of one user from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. The diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ can be expressed as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}=\{\mathbf{d}\in\mathbb{R}^{n}:-r_{i}\leq\mathbf{d}_{i}\leq r_{i},1\leq i\leq n\}$.
4.1. Matrix Mechanism
The matrix mechanism (LMHMR15, ; mckenna2021hdmm, ; xiao2024optimal, ) is a well-established technique for providing differentially private answers to a set of queries. Recently, it has been applied to differentially private online streaming processes (choquette2022multi, ; choquette2024amplified, ), with a particular emphasis on differentially private stochastic gradient descent (DPSGD) (abadi2016deep, ). It factorizes the matrix $\mathbf{W}$ into a strategy matrix $\mathbf{B}$ and a linear transformation matrix $\mathbf{L}$ such that $\mathbf{W}=\mathbf{L}\mathbf{B}$. It first adds independent noise to $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, then uses $\mathbf{L}$ to transform the noisy answers from $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to get noisy answers to $\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, specifically,
(1) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{L}(\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}.$
where $\mathbf{z}$ is drawn from a Gaussian distribution $N(0,\mathbf{I}_{n})$, here the covariance matrix $\mathbf{I}_{n}$ is the identity matrix of size $n\times n$.
We aim to minimize the total variance of noise added, for the mechanism defined in Eq. (1). Because ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}$, $\mathbf{z}\sim N(0,\mathbf{I}_{n})$, the total variance of the noise is $\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$. The matrix mechanism (LMHMR15, ) calculates the matrix $\mathbf{L}$ as $\mathbf{L}=\mathbf{W}\mathbf{B}^{\dagger}$ , where $\mathbf{B}^{\dagger}$ is the Moore–Penrose inverse of $\mathbf{B}$. Here it’s assumed that $\mathbf{B}$ has linearly independent columns so that $\mathbf{B}^{T}\mathbf{B}$ is invertible. Therefore, the total variance can be represented as,
$\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$ $=\operatorname{trace}(\mathbf{W}\mathbf{B}^{\dagger}(\mathbf{W}\mathbf{B}^{\dagger})^{T})=\operatorname{trace}(\mathbf{W}^{T}\mathbf{W}(\mathbf{B}^{T}\mathbf{B})^{-1}).$
We next analyze the sensitivity of the Mechanism 1. Let the function $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, from Theorem 2.5, to satisfy $\rho$-zCDP, it’s required that the sensitivity is bounded $\Delta_{2}(f)\leq\sqrt{2\rho}$. From the definition of sensitivity in Definition 2.4, and the definition of neighboring dataset in Definition 4.2, the sensitivity of Mechanism 1 is
$\Delta_{2}(f)=\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\mathbf{B}\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}\|_{2}=\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}}\|\mathbf{B}\mathbf{d}\|_{2}\leq\sqrt{2\rho}.$
Notice that $\|\mathbf{B}\mathbf{d}\|_{2}$ is a convex and continuous function with respect to $\mathbf{d}$, and the constraint set ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}$ is convex and compact with respect to $\mathbf{d}$. So according to Bauer’s maximum principle (kruvzik2000bauer, ), the function $\|\mathbf{B}\mathbf{d}\|_{2}$ attains its maximum at some extreme point of the set ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}$. Recall that ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}=\{\mathbf{d}\in\mathbb{R}^{n}:-r_{i}\leq\mathbf{d}_{i}\leq r_{i},1\leq i\leq n\}$, so the set of extreme points of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}$ is
(2) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}=\{\mathbf{d}\in\mathbb{R}^{n}:d_{i}\in\{-r_{i},r_{i}\},1\leq i\leq n\}.$
Then the constraint that the sensitivity of Mechanism 1 is at most $\sqrt{2\rho}$ can be written as,
$\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}}\mathbf{d}^{T}\mathbf{B}^{T}\mathbf{B}\mathbf{d}=\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}}\mathbf{d}^{T}\mathbf{B}^{T}\mathbf{B}\mathbf{d}\leq 2\rho.$
Using the trick in (yuan2015optimizing, ), we change the variable $\mathbf{P}=\mathbf{B}^{T}\mathbf{B}$, then the matrix mechanism problem becomes,
(3) $\min_{\mathbf{P}\succ\mathbf{0}}$ $\operatorname{trace}\left(\mathbf{W}^{\top}\mathbf{W}\mathbf{P}^{-1}\right)$
s.t. $\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}}\mathbf{d}^{\top}\mathbf{P}\mathbf{d}\leq 2\rho$
The expression $\mathbf{P}\succ\mathbf{0}$ indicates that the matrix $\mathbf{P}$ is positive definite.
4.2. Existing Methods Don’t Scale
When $n$ is relatively small, it’s possible to enumerate the set ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}$, and the paper (choquette2022multi, ) gives an algorithm to solve the Problem (3). We use the code provided in (choquette2022multi, ) to get the optimal solution $\mathbf{P}^{\star}$ of Problem (3) when $n$ is small, in the hope that we can find some structures in the optimal solution. Note that there are $2^{n}$ vectors in ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}$, when $n\geq 16$, the code runs out of memory on a Google Colab server with a T4 GPU. It is necessary to get an efficient method to address the scalability issue.
To find a structure in the optimal solution $\mathbf{P}^{\star}$ to Problem (3), we numerically calculate the $\mathbf{P}^{\star}$ using the solver in (choquette2022multi, ) under its default parameters. For each n, we randomly generate the per-day constraint $r_{i}\in[1,10]$ and set $\rho=\frac{1}{2}$. Surprisingly, we find that $\mathbf{P}^{\star}$ tends to be a diagonal matrix. Table 7 shows the absolute value of the off-diagonal elements in $\mathbf{P}^{\star}$, the maximum absolute value of $\mathbf{P}^{\star}$ is close to 0. Under the assumption that the optimal solution $\mathbf{P}^{\star}$ is a diagonal matrix, Problem (3) can be reduced to a one-dimensional optimization problem, which can be efficiently solved when $n$ is large.
Table 7. Check the Absolute Value of the Off-Diagonal Elements.
$n$ $\max_{i\neq j}|\mathbf{P}^{\star}[i,j]|$ $\min_{i\neq j}|\mathbf{P}^{\star}[i,j]|$ $\operatorname{mean}_{i\neq j}|\mathbf{P}^{\star}[i,j]|$
2 $6.67*10^{-5}$ $6.67*10^{-5}$ $6.67*10^{-5}$
4 $1.59*10^{-4}$ $1.14*10^{-5}$ $5.38*10^{-5}$
6 $7.26*10^{-5}$ $8.03*10^{-7}$ $1.82*10^{-5}$
8 $4.81*10^{-5}$ $4.70*10^{-8}$ $5.58*10^{-6}$
10 $4.39*10^{-5}$ $1.69*10^{-8}$ $3.03*10^{-6}$
12 $4.36*10^{-5}$ $9.20*10^{-9}$ $2.14*10^{-6}$
14 $4.48*10^{-5}$ $2.79*10^{-10}$ $1.71*10^{-6}$
4.3. Mechanism Design
Following the conjecture that the optimal solution to Problem 3 is a diagonal matrix, we incorporate independent noises into a scaled version of the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. Suppose the length of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ is $n$, let the matrix representation of the scale weight be $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$, here $\frac{1}{\sigma_{i}}>0$ is the scale for ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]$. Let $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$, then $\mathbf{W}=\mathbf{L}\mathbf{B}$.
Recall from Mechanism 1, we first add independent noise to $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, then use $\mathbf{L}$ to transform the noisy answers from $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to get noisy answers to $\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, specifically,
(4) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ $=\mathbf{L}(\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})=\mathbf{L}(diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}}){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})$
$=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z},$
where $\mathbf{z}$ is drawn from a Gaussian distribution $N(0,\mathbf{I}_{n})$, here the covariance matrix $\mathbf{I}_{n}$ is the identity matrix of size $n\times n$.
@@ 4.3. Mechanism Design
The privacy budget is determined by the sensitivity of the function $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, or more precisely, how much the outputs of this mapping can deviate (in terms of $l_{2}$ norm) when transitioning from the input stream ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to a neighboring stream $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$, under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$. The following theorem shows how to calculate the sensitivity of the function $f$ with details provided in Appendix 8.1.
Theorem 4.3 (Sesitivity of Ads-BPC).
@@
The sensitivity of the mechanism 4 under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ is

### v2

@@ 4.2 Mechanism Privacy Budget
Given the constraint that each user can contribute at most $r_{i}$ conversions on day $i$ ($1\leq i\leq n$), the user-level neighboring dataset with bounded per-day contributions, denoted as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$, defines a pair of datasets $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}^{\prime})$ as belonging to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ if ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}^{\prime}$ can be obtained from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by substituting all records (with a maximum of $r_{i}$ records on day $i$) of one user from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. The diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ can be expressed as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}=\{\mathbf{d}\in\mathbb{R}^{n}:-r_{i}\leq\mathbf{d}_{i}\leq r_{i},1\leq i\leq n\}$.
@@
The privacy budget is determined by the sensitivity of the function $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, or more precisely, how much the outputs of this mapping can deviate (in terms of $L_{2}$ norm) when transitioning from the input data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to a neighboring data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}^{\prime}$, under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$. The following theorem shows how to calculate the sensitivity of the function $f$.
Theorem 4.2 (Sensitivity of Mechanism 1).
@@
The sensitivity of the mechanism 1 under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ is